OpenStudy (anonymous):

The measures of three angles of a quadrilateral are 80, 90, and 103 degrees. Find the measure of the fourth angle.

jimthompson5910 (jim_thompson5910):

With any quadrilateral, the angles ALWAYS add up to 360 degrees So simply solve 80+90+103+x = 360 for x to find your answer

OpenStudy (anonymous):

in other words 80+90+103= 273 then 360-273= 87 so your angle is 87.

OpenStudy (anonymous):

cool thanks for the answer

OpenStudy (anonymous):

no problem